To construct a 96% confidence interval for variance we took a sample of 20 objects. The lower and upper critical values are___and___ , respectively.
Answer #1

100(1-\alpha)% = 96% \Rightarrow \alpha = 0.04100(1-\alpha) = 96 \Rightarrow \alpha = 0.04

n=20

\chi^{2}_{n-1, 1-\frac{\alpha}{2}} = \chi^{2}_{19,0.98} = 8.56703542

\chi^{2}_{n-1, \frac{\alpha}{2}} = \chi^{2}_{19,0.02} = 33.68742507

Thus, the critical values are 8.567035 and 33.687425 respectively.